A man start from a point P and walk a distance 5km due north-east to a point Q.From Q he then walked 10km due south-east to a point R. Calculate the shortest distance between P and R.
1
Expert's answer
2021-07-19T09:48:49-0400
"X=5\cos{45}+10\cos{45}\\Y=-5\sin{45}+10\sin{45}\\D=\sqrt{(5\cos{45}+10\cos{45})^2+(-5\sin{45}+10\sin{45})^2}\\D=11.2\ km"